Paul and Jasdeep are going out with their friends Sue and Gertrude to play mini putt. They are going to play in teams. What is the probability that Paul and Jasdeep are partners in the first game? I get 1/6, but the answer is 1/3, pls explain?

Answer by stanbon(75887)

Paul and Jasdeep are going out with their friends Sue and Gertrude to play mini putt. They are going to play in teams. What is the probability that Paul and Jasdeep are partners in the first game? I get 1/6, but the answer is 1/3,
--------------
# of pairs = 4C2 = (4*3)/(1*2) = 2*3 = 6
----------
P(Paul and Jas) = 1/6
--------
You are correct.
Cheers,
Stan H.